OpenStudy (anonymous):

Give an example of a linear transformation T: V -> V which shows that it can happen that T(V) intersect n(T) does not equal zero (is not null).

OpenStudy (anonymous):

what math is this? lol.

OpenStudy (anonymous):

Haha, linear algebra - theory based. They don't have tutors for the class at OSU. I don't know where to take my questions

OpenStudy (anonymous):

oh i see. if you don't get a reply on here then just search the basics on google, it may help you out. i never took this class tbh. or else i wouldve helped you out. :)

OpenStudy (anonymous):

Shoot, yeah. I tried google, and like always it rarely helps with this class. The professor wants proofs, and rarely do people ever type up proofs for these random questions. These weekly problem sets are like death. Exactly like death. Haha, but thanks for your help. I'll go and sulk somewhere else

OpenStudy (anonymous):

omg i hate proofs! i would literally die ! haha.

OpenStudy (anonymous):

lol, are you a math major?

OpenStudy (anonymous):

nope. just here to find my answers ! thats what this site is best for.

OpenStudy (helder_edwin):

do u mean: find \(T:V\to V\) such that \(T(V)\cap\ker T\varsupsetneq\{0\}\)? i mean: that this intersection contains other vector(s) than the zero vector?

OpenStudy (helder_edwin):

@krishnapatel ??

OpenStudy (anonymous):

yes! sorry, our book doesn't use (kernel?) but that is exactly what the problem is asking for

OpenStudy (helder_edwin):

give a few minutes. ok?

OpenStudy (anonymous):

sure, thanks!

OpenStudy (helder_edwin):

r u still there?

OpenStudy (helder_edwin):

sorry. it took a bit long for me to figure this out. do u know derivatives?

OpenStudy (anonymous):

no its fine, yes i know of derivatives and also that derivatives are a linear transformation

OpenStudy (helder_edwin):

great

OpenStudy (helder_edwin):

does the space V need be finite-dimensional?

OpenStudy (anonymous):

From my understanding, yes. But I may be wrong and the question might be looking for such an answer that requires V to be of infinite dimension. A relevant theorem (called the Fundamental Theorem( states: dim T(V) + dim n(T) = dim V. I was thinking we might have to use this somewhere

OpenStudy (helder_edwin):

there is no need for this theorem

OpenStudy (helder_edwin):

well let's consider \(V=\{p(x)\in \mathbb{R}[x]:\deg p(x)\leq n\}\) if V has to be finite-dimensional. if V has to be infinite-dimensional u just put \(V=\mathbb{R}[x]\). then let \(D:V\to V\) be defined by \(D(p(x))=p'(x)\)

OpenStudy (helder_edwin):

got it?

OpenStudy (anonymous):

I understand the transformation, but wouldn't the transformation be from a n dimensional space into a n-1 dimensional space? (Don't you lose a dimension when you take the derivative of a polynomial?)

OpenStudy (helder_edwin):

i a way, u r right. but the problem doesn't say the tranformation has to be onto, i.e. T(V)=V.

OpenStudy (anonymous):

Hmm, but then what does D: V ->V represent?

OpenStudy (anonymous):

instead of say V -> W.

OpenStudy (helder_edwin):

just a linear transformation with domain and codomain equal

OpenStudy (anonymous):

Ah, okay that makes sense. But then what is T(V) or n(T)? would n(T) be a P(n) where n = 0 (Polynomial with degree zero = a constant)?

OpenStudy (helder_edwin):

ok. so \[ \large \ker D=\{p(x)=\alpha:\alpha\in\mathbb{R}\} \] i.e., the kernel is composed by the constant polynomials. ok?

OpenStudy (anonymous):

yes!

OpenStudy (helder_edwin):

great. now the polynomials of the form \(p(x)=kx\) have image \[ \large D(p(x))=D(kx)=(kx)'=k \] a constant polynomial. ok?

OpenStudy (anonymous):

yep!

OpenStudy (helder_edwin):

so we did it!

OpenStudy (helder_edwin):

for this linear transformation \[ \large D(V)\cap\ker D=\ker D\varsupsetneq0 \]

OpenStudy (anonymous):

Sorry that I'm taking a little longer to grasp this. I thought we said that D(V) consists of polynomials of form kx, where as ker D consists of polynomials of form k(x^0). then how is the intersection of those two polynomials ker D?

OpenStudy (anonymous):

Ohhh, is it for all x such that x = 1?

OpenStudy (helder_edwin):

no. \(x\) is an indetermminate. so it is FOR ALL x.

OpenStudy (anonymous):

Yes, actually I think I understand. Correct me if I'm wrong. We have ker D to be a subset of D(V) where the intersection of ker D and D(V) is the set ker D itself.

OpenStudy (anonymous):

which consists of the polynomial functions of the form k(x^0)

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

wow that's beautiful. thank you so much for your help!

OpenStudy (helder_edwin):

u r welcome